A cylindrical tank 2 meters in diameter, 3 meters high os full of oil. If the specific gravity of 

oil is 0.9, what is the mass of oil in the tank


1
Expert's answer
2021-02-15T17:03:18-0500
"m_{oil}=\rho_{oil}V_{cylinder},""m_{oil}=\rho_{w}SG_{oil}\pi r^2h,""m_{oil}=1000\ \dfrac{kg}{m^3}\cdot0.9\cdot\pi\cdot(1\ m)^2\cdot3\ m=8478\ kg."